# AgileRate: Bringing Adaptivity and Robustness to DeFi Lending Markets

### v3

@@ A.1. Alternate objective for demand/supply balancing
A.2. Optimizing interest rate to maximize revenue
@@
Theorem A.1.
Given the user behavior models described in Equations 12 and 11, the revenue maximization problem, as described in Equation 18, can be formulated as follows:
@@
A.3. Formal abstraction of the learning-based controller
@@
A.4. Utility Functions for Strategic Withholding
@@
A.5. Risk Metrics for Default and Liquidation
@@
A.6. M-estimate function for robust RLS
@@
Appendix B Complimentary Material of Evaluation
B.1. Parameter estimation of Compound demand and supply curves
@@ B.3. Robust RLS experiment
These experiments were conducted using synthetic demand and supply functions, following the experimental setup described in Appendix B.2.
B.4. Forgetting factor and estimation error
@@ B.6. Risk controller
Appendix C Proofs
C.1. Proof of theorem 4.1
@@ Appendix C Proofs
If the borrower follows the honest strategy and reports $r^{b}$, during the exploration phase, they will borrow whenever $r\in[r_{\text{min}},r^{b}]$ and pay an average interest rate of $\frac{r_{\text{min}}+r^{b}}{2}$ during those timeslots. There are $\tau\zeta\frac{r^{b}-r_{\text{min}}}{r_{\text{max}}-r_{\text{min}}}$ such timeslots based on the uniform randomization of $\mathcal{C}_{D}$ described in Appendix [A.3]. During the rest of the exploration phase, i.e., when $r>r^{b}$ or during the non-randomized times, we denote the cost of the borrower by $C_{\tau,\zeta}^{r^{b}}$. Finally, during the exploitation phase $T-\tau$, the borrower pays at most rate $r^{b}$ (potentially even lower depending on the rest of the parameters and objective utilization $U^{*}$).
@@
Appendix D Least square estimation related work

### v4

@@ B.3. Alternate objective for demand/supply balancing
B.4. Optimizing interest rate to maximize revenue
@@
Theorem B.1.
Given the user behavior models described in Equations 9 and 8, the revenue maximization problem, as described in Equation 18, can be formulated as follows:
@@
B.5. Formal abstraction of the learning-based controller
@@
B.6. Utility Functions for Strategic Withholding
@@
B.7. Risk Metrics for Default and Liquidation
@@
B.8. M-estimate function for robust RLS
@@
Appendix C Complimentary Material of Evaluation
C.1. Parameter estimation of Compound demand and supply curves
@@ C.3. Robust RLS experiment
These experiments were conducted using synthetic demand and supply functions, following the experimental setup described in Appendix C.2.
C.4. Forgetting factor and estimation error
@@ C.6. Risk controller
Appendix D Proofs
D.1. Proof of theorem 3.1
@@ Appendix D Proofs
If the borrower follows the honest strategy and reports $r^{b}$, during the exploration phase, they will borrow whenever $r\in[r_{\text{min}},r^{b}]$ and pay an average interest rate of $\frac{r_{\text{min}}+r^{b}}{2}$ during those timeslots. There are $\tau\zeta\frac{r^{b}-r_{\text{min}}}{r_{\text{max}}-r_{\text{min}}}$ such timeslots based on the uniform randomization of $\mathcal{C}_{D}$ described in Appendix [B.5]. During the rest of the exploration phase, i.e., when $r>r^{b}$ or during the non-randomized times, we denote the cost of the borrower by $C_{\tau,\zeta}^{r^{b}}$. Finally, during the exploitation phase $T-\tau$, the borrower pays at most rate $r^{b}$ (potentially even lower depending on the rest of the parameters and objective utilization $U^{*}$).
@@
Appendix E Least square estimation related work